OpenStudy (anonymous):

the perimeter of a triangle is 47 centimeters. The longest side is 7 centimeters less than the sum of the other two sides. Twice the shortest side is 4 centimeters less than the longest side. Find the lengths of all sides.

OpenStudy (anonymous):

8,19,20

OpenStudy (anonymous):

great but can you explain to me how you came up with that answer?

OpenStudy (anonymous):

2x+2y-7=47 2x+4 = x+y-7 - x= y-11 etc...